OpenStudy (anonymous):

the length of a rectangle is 10 more than the side of a square, and the width if the rectangle is 10 less than the side of the square. If the area of the square is 125, what is the area of the rectangle?

OpenStudy (anonymous):

and the width of the rectangle is 10 less than the side of the square.

OpenStudy (anonymous):

what is the formula for the area of a square?

OpenStudy (anonymous):

s squared

OpenStudy (anonymous):

I think that one side would be the square root of 125

OpenStudy (anonymous):

the square roots of 125 squared is 125

OpenStudy (anonymous):

very good, although you must say what s means,. s, in this case, means the side of a square..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

right, very good. That means the side of the square is the square root of 125.

OpenStudy (anonymous):

so the length of the rectangle should be the square root of 125 plus 10

OpenStudy (anonymous):

yup, and what about the width?

OpenStudy (anonymous):

that equals 21.18033989

OpenStudy (anonymous):

the width

OpenStudy (anonymous):

1.180339887?

OpenStudy (anonymous):

no no. you dont need to know the value 21.18033989, because I know some magic is going to happen.

OpenStudy (anonymous):

I got 25.74

OpenStudy (anonymous):

Okay. You are correct, but I know of a better way to solve this problem. Are you with me?

OpenStudy (anonymous):

So I would round it to 25

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[length = \sqrt{125} + 10\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[width = \sqrt{125} -10\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

area of a rectangle is length times width. so,\[area = (\sqrt{125}+10) \times (\sqrt{125}-10)\]

OpenStudy (anonymous):

125-100?

OpenStudy (anonymous):

now, there is a little formula which says \[(a+b) \times (a-b) = a ^{2} - b ^{2}\]

OpenStudy (anonymous):

excellent. You are right. see how didnt need to use a calculator?

OpenStudy (anonymous):

the different of two squares?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Cool